# Highest-weight vectors and three-point functions in GKO coset decomposition

### v1

@@ 2.6.2 Degenerate representations
Assume now that $k$ is generic and $\nu\in\mathbb{Z}_{\geq 0}$. Then the representation $\mathcal{M}_{\nu,k}$ has singular vector $f_{0}^{\lambda+1}v_{\nu,k}$. Denote by $\mathcal{L}_{\nu,k}$ the irreducible quotient. Since $k$ is generic we have $\mathcal{L}_{\nu,k}\simeq\operatorname{Ind}^{\widehat{\mathfrak{sl}}(2)}_{\mathfrak{sl}(2)[[t]]\oplus\mathbb{C}K}\mathcal{L}_{\nu}$ where $\mathcal{L}_{\nu}$ is $\nu+1$ dimensional representation of $\mathfrak{sl}(2)$.
Similarly to the construction above we can define the modules $\mathcal{L}_{\nu,k}(x,z)$. Contrary to generic case above the modules $\mathcal{L}_{\nu,k}(x,z)$ do not depend on $x$ as $\widehat{\mathfrak{sl}}(2)_{z}$ modules (up to isomorphism). The analogue of Proposition 2.21 now states that there is a unique up to constant homomorphism of $\mathfrak{sl}(2)\otimes\mathbb{C}[\mathbb{P}^{1}\setminus\{\vec{z}\}]$ modules
@@
The summands satisfy following commutation relations
@@
Finally the fusion rules are equivalent to the following proposition
@@
Let $\nu\in\mathbb{Z}_{\geq 0}$ and $\lambda$ be such that pair $(\lambda,k)$ is generic. Then for any $r\in\{0,1,\dots\nu\}$ there exist and unique up to constant vertex operator $X_{\nu}(x,z):\mathcal{M}_{\lambda-\nu+2r,k}\rightarrow\mathcal{M}_{\lambda,k}$.
@@ 2.6.3 Integrable representations
The similar definitions work for integrable representations of $\widehat{\mathfrak{sl}}(2)_{1}$. Similarly to the above, we can define the representations $\mathcal{L}_{i,1}(x,z)$ of $\widehat{\mathfrak{sl}}(2)_{z}$. These representations do not depend on $x$ up to isomorphism, so they are usually excluded in the construction of vertex operators, but we prefer to keep them.
@@
$\mathtt{m}_{1}:\mathcal{L}_{i,1}(x_{1},z_{1})\otimes\mathcal{L}_{r,1}(x_{2},z_{2})\otimes\mathcal{L}_{j,1}(x_{3},z_{3})\rightarrow\mathbb{C},$ (2.48)
if and only if $r+i+j\in 2\mathbb{Z}$. Composing with isomorphisms $\alpha_{x,z}$ we get a map
$\mathtt{m}_{1}(\vec{x},\vec{z})\colon\mathcal{L}_{i_{1},1}\otimes\mathcal{L}_{i_{2},1}\otimes\mathcal{L}_{i_{3},1}\rightarrow\mathbb{C}.$ (2.49)
This allows to define $Y_{1}(z):\mathcal{L}_{r,1}\rightarrow\operatorname{Hom}(\mathcal{L}_{i,1},\mathcal{L}_{j,1})$. It is clear that $Y_{1}(v_{0}|z)=\operatorname{Id}_{\mathcal{L}_{i,1}}$. Let us define
$b(x,z)=Y_{1}(v_{\frac{1}{2}}|x,z),\quad b(x,z)=b_{0}(z)+xb_{1}(z).$ (2.50)
@@
Recall that $\frac{h_{0}}{2}=\frac{a_{0}}{\sqrt{2}}$.
@@ 2.6.4 Bosonization of vertex operators
where $S(t)=:\!\exp\left(\frac{-2}{\sqrt{2\kappa}}\varphi(t)\right)\!:\beta(t)$ is screening field.
@@
Commutation relation between operators $\mathcal{O}^{(N)}_{r;\nu}(z)$ and $\widehat{\mathfrak{sl}}(2)$ generators have the form
@@
Due to uniqueness of $\mathcal{V}_{\nu}$ it is sufficient to check that right side of satisfies commutation relations (2.39), This is direct computation using Proposition 2.26. ∎
@@ 2.7 Virasoro vertex operators
In the Virasoro case we will follow the same pattern. We keep the notation $t$ for some fixed global coordinate on $\mathbb{P}^{1}$. For any $z\in\mathbb{P}^{1}$ let $\mathrm{Vir}_{z}$ denotes Lie algebra with basis $(t-z)^{n}\partial_{t}$ and central element $C$ with the commutator defined by
@@
which map any element to the direct sum of series expansions at points $z_{i}$. The completions $\overline{\mathrm{Vir}_{z}}$ denote Lie algebra with elements of the form $f(t-z)\partial_{t}+\alpha C$, where $f\in\mathbb{C}[[t-z,(t-z)^{-1}]$ is Laurent polynomials. Clearly $\overline{\mathrm{Vir}_{z}}$ acts on $\mathbb{M}_{P,b}(z)$.
For generic $P_{1},P_{2},P_{3}$ there exists a unique up to constant homomorphism of $\mathrm{Vect}(\mathbb{P}^{1})\setminus\{z_{1},z_{2},z_{3}\})$ modules
@@
For generic $P_{3}$ this map leads to an operator state correspondance map $\mathtt{Y}^{\mathrm{Vir}}(z):\mathbb{M}_{P_{2},b}\rightarrow\operatorname{Hom}_{\mathbb{C}}(\mathbb{M}_{P_{1},b},\overline{\mathbb{M}_{P_{3},b}})$. Let $\Phi_{\Delta}(z)=\mathtt{Y}^{\mathrm{Vir}}(v_{P,b}|z)$, here $\Delta=\Delta(P,b)$. The $\mathrm{Vect}(\mathbb{P}^{1})\setminus\{z_{1},z_{2},z_{3}\})$ invariance of $\mathtt{m}^{\mathrm{Vir}}$ leads to the following commutation relations.
@@
For the $\Delta=\Delta_{m,n}$ Verma module becomes reducible (see Theorem 2.8. We can replace Verma modules by their irreducible quotients $\mathbb{L}_{P_{m,n},b}$. The corresponding vertex operators will be denoted by $\Phi_{m,n}(z)$. The singular vector condition on $v_{P_{m,n},b}$ corresponds to the equation on the vertex operator. In the important case $\Phi_{1,2}(z)$ and $\Phi_{2,1}(z)$ we have (cf. Example 2.9)
@@ 3.1 Definition, decomposition
Let us consider the algebra $\widehat{\mathfrak{sl}}(2)_{1}\oplus\widehat{\mathfrak{sl}}(2)_{k}$ and its module $\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}$. There is an action of diagonal $\widehat{\mathfrak{sl}}(2)_{k+1}$ on this. Let us denote their generators $e_{n}^{\Delta},f_{n}^{\Delta},h_{n}^{\Delta}$.
Note that there are three different actions of the Virasoro algebra on $\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}$ which go from the Sugawara construction. We will call the corresponding currents $L^{(1)}(z),L^{(2)}(z)$ and $L^{\Delta}(z)$.
@@
Let us use the notation $\mathrm{Vir}^{\mathrm{coset}}$ for the algebra generated by $L^{GKO}_{n}$.
@@
For generic $\lambda,k$ there are the following decompositions of $\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}$ as $\mathrm{Vir}^{\mathrm{coset}}\oplus\widehat{\mathfrak{sl}}(2)_{k+1}$ module
@@
$\langle x_{1}\otimes x_{2},y_{1}\otimes y_{2}\rangle=\langle x_{1}|y_{1}\rangle\langle x_{2}|y_{2}\rangle.$ (3.3)
@@ 3.2 Formula for the highest weight vectors
Recall (see sec. 2.4) notation $v_{l}$ for extremal vectors in $L_{i,1}$. For the tensor products of extremal and highest weight vectors we will use notation
@@
Recall that for general $\lambda,k$ Verma module $\mathcal{M}_{\lambda,k}$ is isomorphic to Wakimoto module $\mathcal{F}_{\lambda,k}$ (see Proposition 2.20). Our first goal is to find explicit formulas for vectors $v_{l}(\lambda)$, or actually to their images in $\mathcal{L}_{i,1}\otimes\mathcal{F}_{\lambda,k}$. Let us define current $g(z)$ acting on $\mathcal{L}_{i,1}\otimes\mathcal{F}_{\lambda,k}$
@@
If $2l\in\mathbb{Z}_{\leq 0}$ then $l$-th the highest weight vector in decompositions (3.1a), (3.1b) with respect to $\mathrm{Vir}^{\mathrm{coset}}\oplus\widehat{\mathfrak{sl}}(2)_{k+1}$ is given by
$u_{l}(\lambda)=e^{-g_{0}}v_{l}(\lambda).$ (3.10)
@@
In the theorem above we had $l\leq 0$, let us comment on the other case $l>0$. There is an automorphism $\tau$ of algebra $\widehat{\mathfrak{sl}}(2)_{k}$
@@
where $x\in\widehat{\mathfrak{sl}}(2)_{k}$, $v\in\mathcal{M}_{\lambda,k}$ and $v_{\lambda,k}$ denotes the highest weight vector in $\mathcal{M}_{\lambda,k}$. There exist similar map $\tau\colon\mathcal{L}_{i,1}\rightarrow\mathcal{L}_{1-i,1}$ and for $i\in\{0,1\}$.
@@
Another analog of formula (3.10) for $l>0$ uses conugation an is given in Remark 3.24.
We will first prove Theorem 3.7 by direct computation.
@@
Conjugation by $e^{g_{0}}$ acts on the operators of diagonal $\widehat{\mathfrak{sl}}(2)_{k+1}$ as follows
$e^{g_{0}}e^{\Delta}(w)e^{-g_{0}}=\beta(w),$ (3.15)
$e^{g_{0}}f^{\Delta}(w)e^{-g_{0}}=f^{\Delta}(w)+h^{(1)}(w)\gamma(w)+\gamma^{\prime}(w).$ (3.16)
@@
There are formulas for exponent of adjoint action
$e^{g_{0}}e^{\Delta}(w)e^{-g_{0}}$ $=e^{\Delta}(w)+[g_{0},e^{\Delta}(w)]+\frac{1}{2}[g_{0},[g_{0},e^{\Delta}(w)]]+\dots$ (3.17)
$e^{g_{0}}f^{\Delta}(w)e^{-g_{0}}$ $=f^{\Delta}(w)+[g_{0},f^{\Delta}(w)]+\frac{1}{2}[g_{0},[g_{0},f^{\Delta}(w)]]+\dots$ (3.18)
It means that sufficient to find all commutators. In order to do this we will use OPE (see [FBZ04, Ch 3]),
@@
$e^{g_{0}}e^{\Delta}(w)e^{-g_{0}}=e^{(2)}(w)=\beta(w).$ (3.21)
@@
We decompose the proof into $3$ steps.
Step 1. Let us show that $e^{-g_{0}}v_{-l}(\lambda)$ is the highest weight vector with respect to diagonal $\widehat{\mathfrak{sl}}(2)^{\Delta}$. It is sufficient to consider action $e^{\Delta}_{0}$ and $f^{\Delta}_{1}$. Using Lemma 3.10 we have
$e^{\Delta}_{0}e^{-g_{0}}v_{-l}(\lambda)=e^{-g_{0}}e^{g_{0}}e^{\Delta}_{0}e^{-g_{0}}v_{-l}(\lambda)=e^{-g_{0}}\beta_{0}v_{-l}(\lambda)=0,$ (3.26)
$f^{\Delta}_{1}e^{-g_{0}}v_{-l}(\lambda)=e^{-g_{0}}e^{g_{0}}f^{\Delta}_{1}e^{-g_{0}}v_{-l}(\lambda)=e^{-g_{0}}(f^{(2)}_{1}+f^{(1)}_{1}-\sum_{s+r=1}h^{(1)}_{r}\gamma_{s}-\gamma_{1})v_{-l}(\lambda)=0.$ (3.27)
Step 2. It follows from the decompositions (3.1) that there exist only one up to proportionality vector $v\in W_{i}$ which is highest weight vector for $\widehat{\mathfrak{sl}}(2)^{\Delta}$ and the same eigenvalues of $h_{0}^{\Delta}v$ as for $u_{l}(\lambda)$. and $L_{0}^{GKO}$. Hence, the vectors $u_{l}(\lambda)$ and $e^{-g_{0}}v_{l}(\lambda)$ are proportional.
@@
$\langle e^{-g_{0}}v_{-l}(\lambda)|v_{-l}(\lambda)\rangle=1.$ (3.28)
which is straightforward. ∎
@@
$I(z)=e^{\frac{\varphi(z)}{\sqrt{2\kappa}}}e^{-g_{0}}b_{1}(z)e^{g_{0}}.$ (3.34)
@@
2) For $2l\leq\mathbb{Z}_{\leq 0}$ we have
$u_{l}(\lambda)=e^{-g_{0}}v_{l}(\lambda).$ (3.37)
The exponent $d(l,\lambda,k)$ in (3.36) means that the leading term of $I(z)u_{l}(\lambda)$ coincides with $u_{l-\frac{1}{2}}(\lambda+1)$.
@@
$\oint_{C_{0}}\frac{dz}{z^{d(s,\lambda,k)}}I(z)e^{-g_{0}}v_{l}(\lambda)=(-1)^{l(2l-1)}e^{-g_{0}}v_{l-1/2}(\lambda+1).$ (3.39)
@@
$e^{-g_{0}}v_{l}(\lambda)=\left(\prod\nolimits_{2m\in\mathbb{Z},l<m\leq 0}c_{m,\lambda-2l+2m}\right)u_{l}(\lambda).$ (3.40)
Taking into account correct normalization property (3.28) we conclude that $c_{l}=1$ for $l\leq 0$. This finishes the proof. ∎
Note that $c_{l,\lambda}$ appearing in the proof above is non-trivial for $l>0$. We will compute them below.

### v3

@@ 2.6.2 Degenerate representations
Assume now that $k$ is generic and $\nu\in\mathbb{Z}_{\geq 0}$. Then the representation $\mathcal{M}_{\nu,k}$ has singular vector $f_{0}^{\nu+1}v_{\nu,k}$. Denote by $\mathcal{L}_{\nu,k}$ the irreducible quotient. Since $k$ is generic we have $\mathcal{L}_{\nu,k}\simeq\operatorname{Ind}^{\widehat{\mathfrak{sl}}(2)}_{\mathfrak{sl}(2)[[t]]\oplus\mathbb{C}K}\mathcal{L}_{\nu}$ where $\mathcal{L}_{\nu}$ is $\nu+1$ dimensional representation of $\mathfrak{sl}(2)$.
Similarly to the construction above we can define modules $\mathcal{L}_{\nu,k}(x,z)$. Contrary to generic case above the modules $\mathcal{L}_{\nu,k}(x,z)$ do not depend on $x$ as $\widehat{\mathfrak{sl}}(2)_{z}$ modules (up to isomorphism). The analogue of Proposition 2.21 now states that there is a unique up to constant homomorphism of $\mathfrak{sl}(2)\otimes\mathbb{C}[\mathbb{P}^{1}{\setminus}\{\vec{z}\}]$ modules
@@
The summands satisfy the following commutation relations
@@
Finally the fusion rules are equivalent to the following proposition.
@@
Let $\nu\in\mathbb{Z}_{\geq 0}$ and $\lambda$ be such that pair $(\lambda,k)$ is generic. Then, for any $r\in\{0,1,\dots\nu\}$ there exist a unique (up to constant) vertex operator $X_{\nu}(x,z):\mathcal{M}_{\lambda-\nu+2r,k}\rightarrow\mathcal{M}_{\lambda,k}$.
@@ 2.6.3 Integrable representations
The similar definitions work for integrable representations of $\widehat{\mathfrak{sl}}(2)_{1}$. Similarly to the above, we can define the representations $\mathcal{L}_{i,1}(x,z)$, $(i=0,1)$ of $\widehat{\mathfrak{sl}}(2)_{z}$. These representations do not depend on $x$ up to isomorphism, so these parameters are usually excluded in the construction of vertex operators, but we prefer to keep them.
@@
$\mathtt{m}_{1}:\mathcal{L}_{i,1}(x_{1},z_{1})\otimes\mathcal{L}_{j,1}(x_{2},z_{2})\otimes\mathcal{L}_{r,1}(x_{3},z_{3})\rightarrow\mathbb{C},$ (2.48)
if and only if $i+j+r\in 2\mathbb{Z}$. Composing with isomorphisms $\alpha_{x,z}$ we get a map
$\mathtt{m}_{1}(\vec{x},\vec{z})\colon\mathcal{L}_{i,1}\otimes\mathcal{L}_{j,1}\otimes\mathcal{L}_{r,1}\rightarrow\mathbb{C}.$ (2.49)
This allows to define $Y_{1}(z):\mathcal{L}_{j,1}\rightarrow\operatorname{Hom}(\mathcal{L}_{i,1},\mathcal{L}_{r,1})$. It is clear that $Y_{1}(v_{0}|z)=\operatorname{Id}_{\mathcal{L}_{i,1}}$. Let us define
$b(x,z)=Y_{1}(v_{{1}/{2}}|x,z),\quad b(x,z)=b_{0}(z)+xb_{1}(z).$ (2.50)
@@
Recall that ${h_{0}}={a_{0}}{\sqrt{2}}$.
@@ 2.6.4 Bosonization of vertex operators
where $S(t)=:\!\exp\left(\frac{-2}{\sqrt{2\kappa}}\varphi(t)\right)\!:\beta(t)$ is a screening field.
@@
Commutation relations between operators $\mathcal{O}^{(N)}_{r;\nu}(z)$ and $\widehat{\mathfrak{sl}}(2)$ generators have the form
@@
Due to uniqueness of $\mathcal{V}_{\nu}$ it is sufficient to check that right side of satisfies commutation relations (2.39). This is direct computation using Proposition 2.26. ∎
@@ 2.7 Virasoro vertex operators
In the Virasoro case we follow the same pattern. We keep the notation $t$ for some fixed global coordinate on $\mathbb{P}^{1}$. For any $z\in\mathbb{P}^{1}$ let $\mathrm{Vir}_{z}$ denotes Lie algebra with basis $(t-z)^{n}\partial_{t}$ and central element $C$ with the commutator defined by
@@
which map any element to the direct sum of series expansions at points $z_{i}$. The completions $\overline{\mathrm{Vir}_{z}}$ denote Lie algebra with elements of the form $f(t-z)\partial_{t}+\alpha C$, where $f\in\mathbb{C}[[t-z,(t-z)^{-1}]$ is a Laurent series. Clearly $\overline{\mathrm{Vir}_{z}}$ acts on $\mathbb{M}_{P,b}(z)$.
For generic $P_{1},P_{2},P_{3}$ there exists a unique up to constant homomorphism of $\mathrm{Vect}\big{(}\mathbb{P}^{1}{\setminus}\{z_{1},z_{2},z_{3}\}\big{)}$ modules
@@
For generic $P_{3}$ this map leads to an operator state correspondence map $\mathtt{Y}^{\mathrm{Vir}}(z):\mathbb{M}_{P_{2},b}\rightarrow\operatorname{Hom}_{\mathbb{C}}(\mathbb{M}_{P_{1},b},\overline{\mathbb{M}_{P_{3},b}})$. Let $\Phi_{\Delta}(z)=\mathtt{Y}^{\mathrm{Vir}}(v_{P,b}|z)$, here $\Delta=\Delta(P,b)$. The $\mathrm{Vect}\big{(}\mathbb{P}^{1}{\setminus}\{z_{1},z_{2},z_{3}\}\big{)}$ invariance of $\mathtt{m}^{\mathrm{Vir}}$ leads to the following commutation relations.
@@
For the $\Delta=\Delta_{m,n}$ Verma module becomes reducible (see Theorem 2.8). We can replace Verma modules by their irreducible quotients $\mathbb{L}_{P_{m,n},b}$. The corresponding vertex operators will be denoted by $\Phi_{m,n}(z)$. The singular vector condition on $v_{P_{m,n},b}$ corresponds to the equation on the vertex operator. In the important cases $\Phi_{1,2}(z)$ and $\Phi_{2,1}(z)$ we have (cf. Example 2.9)
@@ 3.1 Definition, decomposition
Let us consider the algebra $\widehat{\mathfrak{sl}}(2)_{1}\oplus\widehat{\mathfrak{sl}}(2)_{k}$ and its module $\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}$. There is an action of diagonal $\widehat{\mathfrak{sl}}(2)_{k+1}$ on this space. Let us denote the generators of this algebra by $e_{n}^{\Delta},f_{n}^{\Delta},h_{n}^{\Delta}$.
Note that there are three different actions of the Virasoro algebra on $\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}$ which come from the Sugawara construction. We will call the corresponding currents by $L^{(1)}(z),L^{(2)}(z)$ and $L^{\Delta}(z)$.
@@
Let us use the notation $\mathrm{Vir}^{\mathrm{coset}}$ for the algebra generated by $L^{GKO}_{n}$. The following result is standard, we follow [BFL16] for the statement and sketch of the proof.
@@
For generic $\lambda,k$ there are the following decompositions of $\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}$ as $\mathrm{Vir}^{\mathrm{coset}}\oplus\widehat{\mathfrak{sl}}(2)_{k+1}$ modules
@@
$\langle x_{1}\otimes x_{2},y_{1}\otimes y_{2}\rangle=\langle x_{1},y_{1}\rangle\langle x_{2},y_{2}\rangle.$ (3.3)
@@ 3.2 Formula for the highest weight vectors
Recall (see sec. 2.4) notation $v_{l}$ for extremal vectors in $\mathcal{L}_{i,1}$. For the tensor products of extremal and highest weight vectors we will use notation
@@
Recall that for general $\lambda,k$ Verma module $\mathcal{M}_{\lambda,k}$ is isomorphic to Wakimoto module $\mathcal{F}_{\lambda,k}$ (see Proposition 2.20). Our first goal is to find explicit formulas for vectors $u_{l}(\lambda)$, or actually to their images in $\mathcal{L}_{i,1}\otimes\mathcal{F}_{\lambda,k}$. Let us define current $g(z)$ acting on $\mathcal{L}_{i,1}\otimes\mathcal{F}_{\lambda,k}$
@@
For generic $\lambda,k$ and $2l\in\mathbb{Z}_{\leq 0}$ then $l$-th highest weight vector in decompositions (3.1) with respect to $\mathrm{Vir}^{\mathrm{coset}}\oplus\widehat{\mathfrak{sl}}(2)_{k+1}$ is given by
$u_{l}(\lambda)={\mathrm{e}}^{-g_{0}}v_{l}(\lambda).$ (3.10)
Remark 3.8.
In the proof below we will see that the right side of (3.10) is defines the highest weight vector in $\mathcal{L}_{i,1}\otimes\mathcal{F}_{\lambda,k}$ for arbitrary $\lambda,k$. On the other hand, in order to obtain highest weight vector in $\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}$ we use Proposition 2.20, i.e. condition that $\lambda,k$ are generic is essential for the proof. Furthermore, the definition of vectors $u_{l}(\lambda)$ was based on decomposition (3.1), i.e. only for generic values of parameters. See also Remark 3.19 below.
@@
In the theorem above we had $l\leq 0$. Let us comment on the other case $l>0$. There is an automorphism $\tau$ of algebra $\widehat{\mathfrak{sl}}(2)_{k}$
@@
where $x\in\widehat{\mathfrak{sl}}(2)_{k}$, $v\in\mathcal{M}_{\lambda,k}$ and $v_{\lambda,k}$ denotes the highest weight vector in $\mathcal{M}_{\lambda,k}$. There exist similar map $\tau\colon\mathcal{L}_{i,1}\rightarrow\mathcal{L}_{1-i,1}$, for $i\in\{0,1\}$.
@@
Another analog of formula (3.10) for $l>0$ uses conjugated operator $g_{0}^{\dagger}$. It is given in Remark 3.25.
We will first prove Theorem 3.7 by a direct computation.
@@
Conjugation by ${\mathrm{e}}^{g_{0}}$ acts on the operators of diagonal $\widehat{\mathfrak{sl}}(2)_{k+1}$ as follows
${\mathrm{e}}^{g_{0}}e^{\Delta}(w){\mathrm{e}}^{-g_{0}}=\beta(w),$ (3.15)
${\mathrm{e}}^{g_{0}}f^{\Delta}(w){\mathrm{e}}^{-g_{0}}=f^{\Delta}(w)+h^{(1)}(w)\gamma(w)+\gamma^{\prime}(w).$ (3.16)
@@
Th exponent of adjoint action has the form
${\mathrm{e}}^{g_{0}}e^{\Delta}(w){\mathrm{e}}^{-g_{0}}$ $=e^{\Delta}(w)+[g_{0},e^{\Delta}(w)]+\frac{1}{2}[g_{0},[g_{0},e^{\Delta}(w)]]+\dots\;,$ (3.17)
${\mathrm{e}}^{g_{0}}f^{\Delta}(w){\mathrm{e}}^{-g_{0}}$ $=f^{\Delta}(w)+[g_{0},f^{\Delta}(w)]+\frac{1}{2}[g_{0},[g_{0},f^{\Delta}(w)]]+\dots\;.$ (3.18)
Hence it is sufficient to find all commutators. In order to do this we will use OPE (see [FBZ04, Ch 3])
@@
${\mathrm{e}}^{g_{0}}e^{\Delta}(w){\mathrm{e}}^{-g_{0}}=e^{(2)}(w)=\beta(w).$ (3.21)
@@
We decompose the proof into three steps.
Step 1. Let us show that ${\mathrm{e}}^{-g_{0}}v_{-l}(\lambda)$ is the highest weight vector with respect to diagonal $\widehat{\mathfrak{sl}}(2)^{\Delta}$. It is sufficient to consider action $e^{\Delta}_{0}$ and $f^{\Delta}_{1}$. Using Lemma 3.11 we have
$e^{\Delta}_{0}{\mathrm{e}}^{-g_{0}}v_{-l}(\lambda)={\mathrm{e}}^{-g_{0}}{\mathrm{e}}^{g_{0}}e^{\Delta}_{0}{\mathrm{e}}^{-g_{0}}v_{-l}(\lambda)={\mathrm{e}}^{-g_{0}}\beta_{0}v_{-l}(\lambda)=0,$ (3.26)
$f^{\Delta}_{1}{\mathrm{e}}^{-g_{0}}v_{-l}(\lambda)={\mathrm{e}}^{-g_{0}}{\mathrm{e}}^{g_{0}}f^{\Delta}_{1}{\mathrm{e}}^{-g_{0}}v_{-l}(\lambda)={\mathrm{e}}^{-g_{0}}\Big{(}f^{(2)}_{1}+f^{(1)}_{1}-\sum_{s+r=1}h^{(1)}_{r}\gamma_{s}-\gamma_{1}\Big{)}v_{-l}(\lambda)=0.$ (3.27)
Step 2. It follows from the decompositions (3.1) that there exist only one up to proportionality vector $v\in\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}$ which is the highest weight vector for $\widehat{\mathfrak{sl}}(2)^{\Delta}$ and has the same eigenvalues of $h_{0}^{\Delta}$ and $L_{0}^{GKO}$ as $u_{l}(\lambda)$. Hence, the vectors $u_{l}(\lambda)$ and ${\mathrm{e}}^{-g_{0}}v_{l}(\lambda)$ are proportional.
@@
$\langle{\mathrm{e}}^{-g_{0}}v_{-l}(\lambda),v_{-l}(\lambda)\rangle=1.$ (3.28)
Note that in the series ${\mathrm{e}}^{-g_{0}}v_{-l}(\lambda)=\sum_{j\geq 0}(-g_{0})^{j}v_{-l}(\lambda)/j!$ all summands except the first one has $h_{0}^{(1)}$ gradation different from the one of $v_{-l}(\lambda)$. Hence they are orthogonal to $v_{-l}(\lambda)$, ∎
@@
$I(z)={\mathrm{e}}^{\frac{\varphi(z)}{\sqrt{2\kappa}}}{\mathrm{e}}^{-g_{0}}b_{1}(z){\mathrm{e}}^{g_{0}}.$ (3.34)
@@
2) For $2l\in\mathbb{Z}_{\leq 0}$ we have
$u_{l}(\lambda)={\mathrm{e}}^{-g_{0}}v_{l}(\lambda).$ (3.37)
The exponent $d(l,\lambda,k)$ in (3.36) means that the leading term of $I(z)u_{l}(\lambda)$ coincides with $u_{l-{1}/{2}}(\lambda+1)$.
@@
$\oint_{C_{0}}\frac{dz}{z^{d(s,\lambda,k)}}I(z){\mathrm{e}}^{-g_{0}}v_{l}(\lambda)=(-1)^{l(2l-1)}{\mathrm{e}}^{-g_{0}}v_{l-1/2}(\lambda+1).$ (3.39)
@@
${\mathrm{e}}^{-g_{0}}v_{l}(\lambda)=\left(\prod\nolimits_{2m\in\mathbb{Z},l<m\leq 0}c_{m,\lambda-2l+2m}\right)u_{l}(\lambda).$ (3.40)
Taking into account correct normalization property (3.28) we conclude that $c_{l\lambda}=1$ for $l\leq 0$. This finishes the proof. ∎
Note that $c_{l,\lambda}$ appearing in the proof above is non-trivial for $l>0$. We will discuss them belowm see Example 4.7.